# Optimizing importance weighting in the presence of sub-population shifts

### v1

@@ B.2 The asymptotic bias and variance of the coefficients
$\mathbb{E}[\hat{\bm{\beta}}(p)|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}}]$ $=\mathbb{E}\left[(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\frac{1}{n}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{y}_{\mathrm{tr}}|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}}\right]$
$=\mathbb{E}\left[\left(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{y}_{1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{y}_{0}\right)|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}}\right]$
$=\left(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\mathbb{E}[\bm{y}_{1}|\bm{X}]+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\mathbb{E}[\bm{y}_{0}|\bm{X}]\right)$
$=\left(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}\bm{\beta}^{1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0}\bm{\beta}^{0}\right).$
@@
$\mathrm{Var}(\hat{\bm{\beta}}(p)|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}})$ $=\mathrm{Var}((\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{y}_{\mathrm{tr}}|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}})$
$=(\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\mathrm{Var}(\bm{y}_{\mathrm{tr}}|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}})\bm{W}\bm{X}_{\mathrm{tr}}(\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}$
$=\sigma^{2}(\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{W}\bm{X}_{\mathrm{tr}}(\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}$ (30)
Similar to the data-generating process in Section 4, we assume here that each $\bm{x}$ follows the same distribution regardless of the group. Given this, we can state $\mathbb{E}[\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}]=\mathbb{E}[\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0}]=\bm{\Sigma}$, e.g. they have the same expectation. Assuming $\bm{\Sigma}$ is a positive definite matrix with finite elements, $\frac{1}{n_{1}}\bm{X}_{1}^{T}\bm{X}_{1},\frac{1}{n_{0}}\bm{X}_{0}^{T}\bm{X}_{0}$ converge in probability to it, e.g.
$\underset{n_{1}\rightarrow\infty}{\mathrm{plim}}\frac{1}{n_{1}}\bm{X}_{1}^{T}\bm{X}_{1}=\underset{n_{0}\rightarrow\infty}{\mathrm{plim}}\frac{1}{n_{0}}\bm{X}_{0}^{T}\bm{X}_{0}=\bm{\Sigma}.$ (31)
@@
$\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}=\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}=\bm{\Sigma}+\mathcal{O}_{p}(\frac{1}{\sqrt{n}}).$ (32)
@@
$\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0}\right)=p\bm{\Sigma}+(1-p)\bm{\Sigma}+\mathcal{O}_{p}(\frac{1}{\sqrt{n}})=\bm{\Sigma}+\mathcal{O}_{p}(\frac{1}{\sqrt{n}}),$
@@
$\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0}\right)^{-1}$ $=\left(\bm{\Sigma}+\mathcal{O}_{p}(\frac{1}{\sqrt{n}})\right)^{-1}$
@@
The $\mathcal{O}_{p}(\frac{1}{\sqrt{n}})$ indicates $\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1},\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}$ are bounded in probability (see Vaart (1998) Section 2.2). Thus, for a large $n$, we can approximate $\mathbb{E}\left[\left(\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\right]$ using $\bm{\Sigma}^{-1}$.
@@
$=\mathbb{E}_{\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}}}\left[\mathbb{E}\left[\left(\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{T}\bm{X}_{\mathrm{tr},1}\bm{\beta}^{1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{T}\bm{X}_{\mathrm{tr},0}\bm{\beta}^{0}\right)\right]\right]$
@@
$\underset{n\rightarrow\infty}{\mathrm{plim}}\frac{1}{n}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{W}\bm{X}_{\mathrm{tr}}=\left(\frac{p^{2}}{p_{\mathrm{tr}}}+\frac{(1-p)^{2}}{(1-p_{\mathrm{tr}})})\bm{\Sigma}\right).$ (35)
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathrm{Var}(\hat{\bm{\beta}}(\bm{w})|D_{\mathrm{tr}})\right]$ $=\mathbb{E}_{\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}}}\left[\sigma^{2}(\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{W}\bm{X}_{\mathrm{tr}}(\bm{X}_{\mathrm{tr}}^{T}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\right]$
@@ B.3 The expected loss
If we substitute in our approximations derived in Equation 36, Equation 34 in Equation B.1, we get
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{p_{\mathrm{te}}(y)}[(y-\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|\bm{x}_{\mathrm{te}}]\right]$ $\approx\left(\sum^{G}_{g=1}p_{\mathrm{te},g}((\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{g}-\bm{x}_{\mathrm{te}}^{T}(p\bm{\beta}^{1}+(1-p)\bm{\beta}^{0}))^{2}+\sigma^{2}_{g}\right)$
$+\bm{x}_{\mathrm{te}}^{T}\frac{1}{n}\sigma^{2}\left(\frac{p^{2}}{p_{\mathrm{tr},1}}+\frac{(1-p)^{2}}{(1-p_{\mathrm{tr},1})}\right)\bm{\Sigma}^{-1}\bm{x}_{\mathrm{te}}.$
@@
$\mathbb{E}[(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}^{g}-\bm{x}_{\mathrm{te}}^{T}\left(p\bm{\beta}^{1}+(1-p)\bm{\beta}^{0}\right)])^{2}]$ $=\mathbb{E}[(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}_{d}^{g})^{2}],$
@@
$(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}_{d}^{g})^{2}$ $=(\sum^{p}_{j=1}x_{j}\beta_{d,j}^{g})(\sum^{p}_{j=1}x_{j}\beta_{d,j}^{g})=\mathrm{Tr}(\bm{x}_{\mathrm{te}}\bm{x}_{\mathrm{te}}^{T}\bm{\beta}_{d}^{g}\bm{\beta}_{d}^{{g}^{T}}),$
$\mathbb{E}[(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}_{d}^{g})^{2}]$ $=\mathrm{Tr}(\mathbb{E}[\bm{x}_{\mathrm{te}}\bm{x}_{\mathrm{te}}^{T}]\mathbb{E}[\bm{\beta}_{d}^{g}\bm{\beta}_{d}^{{g}^{T}}]),$ (Since they are independent)
$\mathbb{E}[\bm{x}_{\mathrm{te}}\bm{x}_{\mathrm{te}}^{T}]$ $=\begin{pmatrix}1&\bm{0}^{T}\\ \bm{0}&\bm{\Sigma}.\end{pmatrix}.$
@@
$\bm{\beta}_{d}^{(1)}\bm{\beta}_{d}^{{(1)}^{T}}$ $=(1-p)^{2}\bm{\beta}_{d}\bm{\beta}_{d}^{T}$
$\bm{\beta}_{d}^{(0)}\bm{\beta}_{d}^{{(0)}^{T}}$ $=p^{2}\bm{\beta}_{d}\bm{\beta}_{d}^{T}.$
@@
$\mathbb{E}[(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}_{d}^{(1)})^{2}]$ $=(1-p)^{2}(a_{1}-a_{0})^{2},$
$\mathbb{E}[(\bm{x}_{\mathrm{te}}^{T}\bm{\beta}_{d}^{(0)})^{2}]$ $=p^{2}(a_{1}-a_{0})^{2}.$
@@
$\mathbb{E}[\bm{x}_{\mathrm{te}}^{T}\frac{\sigma^{2}}{n}(\frac{p^{2}}{p_{\mathrm{tr}}}+\frac{(1-p)^{2}}{(1-p_{\mathrm{tr}})})\bm{\Sigma}^{-1}\bm{x}_{\mathrm{te}}]$ $=\sigma^{2}\left(\frac{p^{2}}{p_{\mathrm{tr}}}+\frac{(1-p)^{2}}{(1-p_{\mathrm{tr}})}\right)\frac{1}{n}\mathbb{E}[\bm{x}_{\mathrm{te}}^{T}\begin{pmatrix}1&\bm{0}^{T}\\ \bm{0}&\frac{1}{\gamma}\bm{I}\end{pmatrix}\bm{x}_{\mathrm{te}}]$
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\!\left[\mathbb{E}_{(y,\bm{x})\sim\mathcal{P}_{\mathrm{te}}}[(y-\bm{x}_{\mathrm{te}}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|D_{\mathrm{tr}}]\right]$ $\approx\mathcal{B}^{2}(p_{\mathrm{te}},p,a_{1},a_{0})+\mathcal{V}(\sigma^{2},p,p_{\mathrm{tr}},n,d)+\sigma^{2}.$ (37)

### v2

@@ B.2 The asymptotic bias and variance of the coefficients
$\mathbb{E}[\hat{\bm{\beta}}(p)|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}}]$ $=\mathbb{E}\left[(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\frac{1}{n}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{y}_{\mathrm{tr}}|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}}\right]$
$=\mathbb{E}\left[\left(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{y}_{1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{y}_{0}\right)|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}}\right]$
$=\left(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\mathbb{E}[\bm{y}_{1}|\bm{X}]+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\mathbb{E}[\bm{y}_{0}|\bm{X}]\right)$
$=\left(\frac{1}{n}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}\bm{\beta}^{1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}\bm{\beta}^{0}\right).$
@@
$\mathrm{Var}(\hat{\bm{\beta}}(p)|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}})$ $=\mathrm{Var}((\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{y}_{\mathrm{tr}}|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}})$
$=(\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\mathrm{Var}(\bm{y}_{\mathrm{tr}}|\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}})\bm{W}\bm{X}_{\mathrm{tr}}(\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}$
$=\sigma^{2}(\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{W}\bm{X}_{\mathrm{tr}}(\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}$ (30)
Similar to the data-generating process in Section 4, we assume here that each $\bm{x}$ follows the same distribution regardless of the group. Given this, we can state $\mathbb{E}[\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}]=\mathbb{E}[\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}]=\bm{\Sigma}$, e.g. they have the same expectation. Assuming $\bm{\Sigma}$ is a positive definite matrix with finite elements, $\frac{1}{n_{1}}\bm{X}_{1}^{\mathrm{T}}\bm{X}_{1},\frac{1}{n_{0}}\bm{X}_{0}^{\mathrm{T}}\bm{X}_{0}$ converge in probability to it, e.g.
$\underset{n_{1}\rightarrow\infty}{\mathrm{plim}}\frac{1}{n_{1}}\bm{X}_{1}^{\mathrm{T}}\bm{X}_{1}=\underset{n_{0}\rightarrow\infty}{\mathrm{plim}}\frac{1}{n_{0}}\bm{X}_{0}^{\mathrm{T}}\bm{X}_{0}=\bm{\Sigma}.$ (31)
@@
$\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}=\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}=\bm{\Sigma}+\mathcal{O}_{p}(\frac{1}{\sqrt{n}}).$ (32)
@@
$\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}\right)=p\bm{\Sigma}+(1-p)\bm{\Sigma}+\mathcal{O}_{p}(\frac{1}{\sqrt{n}})=\bm{\Sigma}+\mathcal{O}_{p}(\frac{1}{\sqrt{n}}),$
@@
$\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}\right)^{-1}$ $=\left(\bm{\Sigma}+\mathcal{O}_{p}(\frac{1}{\sqrt{n}})\right)^{-1}$
@@
The $\mathcal{O}_{p}(\frac{1}{\sqrt{n}})$ indicates $\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1},\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}$ are bounded in probability (see van der Vaart (1998) Section 2.2). Thus, for a large $n$, we can approximate $\mathbb{E}\left[\left(\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\right]$ using $\bm{\Sigma}^{-1}$.
@@
$=\mathbb{E}_{\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}}}\left[\mathbb{E}\left[\left(\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}}\right)^{-1}\left(p\frac{1}{n_{1}}\bm{X}_{\mathrm{tr},1}^{\mathrm{T}}\bm{X}_{\mathrm{tr},1}\bm{\beta}^{1}+(1-p)\frac{1}{n_{0}}\bm{X}_{\mathrm{tr},0}^{\mathrm{T}}\bm{X}_{\mathrm{tr},0}\bm{\beta}^{0}\right)\right]\right]$
@@
$\underset{n\rightarrow\infty}{\mathrm{plim}}\frac{1}{n}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{W}\bm{X}_{\mathrm{tr}}=\left(\frac{p^{2}}{p_{\mathrm{tr}}}+\frac{(1-p)^{2}}{(1-p_{\mathrm{tr}})}\right)\bm{\Sigma}.$ (35)
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathrm{Var}(\hat{\bm{\beta}}(\bm{w})|D_{\mathrm{tr}})\right]$ $=\mathbb{E}_{\bm{g}_{\mathrm{tr}},\bm{X}_{\mathrm{tr}}}\left[\sigma^{2}(\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{W}\bm{X}_{\mathrm{tr}}(\bm{X}_{\mathrm{tr}}^{\mathrm{T}}\bm{W}\bm{X}_{\mathrm{tr}})^{-1}\right]$
@@ B.3 The expected loss
If we substitute in our approximations derived in Equations 34,36 in Equation B.1, we get
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\left[\mathbb{E}_{p_{\mathrm{te}}(y)}[(y-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|\bm{x}_{\mathrm{te}}]\right]$ $\approx\left(\sum^{G}_{g=1}p_{\mathrm{te},g}((\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{g}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}(p\bm{\beta}^{1}+(1-p)\bm{\beta}^{0}))^{2}+\sigma^{2}_{g}\right)$
$+\bm{x}_{\mathrm{te}}^{\mathrm{T}}\frac{1}{n}\sigma^{2}\left(\frac{p^{2}}{p_{\mathrm{tr},1}}+\frac{(1-p)^{2}}{(1-p_{\mathrm{tr},1})}\right)\bm{\Sigma}^{-1}\bm{x}_{\mathrm{te}}.$
@@
$\mathbb{E}\left[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}^{g}-\bm{x}_{\mathrm{te}}^{\mathrm{T}}(p\bm{\beta}^{1}+(1-p)\bm{\beta}^{0})\right)^{2}]$ $=\mathbb{E}[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}_{d}^{g})^{2}],$
@@
$(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}_{d}^{g})^{2}$ $=(\sum^{p}_{j=1}x_{j}\beta_{d,j}^{g})(\sum^{p}_{j=1}x_{j}\beta_{d,j}^{g})=\mathrm{Tr}(\bm{x}_{\mathrm{te}}\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}_{d}^{g}\bm{\beta}_{d}^{{g}^{\mathrm{T}}}),$
$\mathbb{E}[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}_{d}^{g})^{2}]$ $=\mathrm{Tr}(\mathbb{E}[\bm{x}_{\mathrm{te}}\bm{x}_{\mathrm{te}}^{\mathrm{T}}]\mathbb{E}[\bm{\beta}_{d}^{g}\bm{\beta}_{d}^{{g}^{\mathrm{T}}}]),$ (Since they are independent)
$\mathbb{E}[\bm{x}_{\mathrm{te}}\bm{x}_{\mathrm{te}}^{\mathrm{T}}]$ $=\begin{pmatrix}1&\bm{0}^{\mathrm{T}}\\ \bm{0}&\bm{\Sigma}.\end{pmatrix}$.
@@
$\bm{\beta}_{d}^{(1)}\bm{\beta}_{d}^{{(1)}^{\mathrm{T}}}$ $=(1-p)^{2}\bm{\beta}_{d}\bm{\beta}_{d}^{\mathrm{T}}$
$\bm{\beta}_{d}^{(0)}\bm{\beta}_{d}^{{(0)}^{\mathrm{T}}}$ $=p^{2}\bm{\beta}_{d}\bm{\beta}_{d}^{\mathrm{T}}.$
@@
$\mathbb{E}[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}_{d}^{(1)})^{2}]$ $=(1-p)^{2}(a_{1}-a_{0})^{2},$
$\mathbb{E}[(\bm{x}_{\mathrm{te}}^{\mathrm{T}}\bm{\beta}_{d}^{(0)})^{2}]$ $=p^{2}(a_{1}-a_{0})^{2}.$
@@
$\mathbb{E}[\bm{x}_{\mathrm{te}}^{\mathrm{T}}\frac{\sigma^{2}}{n}(\frac{p^{2}}{p_{\mathrm{tr}}}+\frac{(1-p)^{2}}{(1-p_{\mathrm{tr}})})\bm{\Sigma}^{-1}\bm{x}_{\mathrm{te}}]$ $=\sigma^{2}\left(\frac{p^{2}}{p_{\mathrm{tr}}}+\frac{(1-p)^{2}}{(1-p_{\mathrm{tr}})}\right)\frac{1}{n}\mathbb{E}[\bm{x}_{\mathrm{te}}^{\mathrm{T}}\begin{pmatrix}1&\bm{0}^{\mathrm{T}}\\ \bm{0}&\frac{1}{\gamma}\bm{I}\end{pmatrix}\bm{x}_{\mathrm{te}}]$
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\!\left[\mathbb{E}_{(y,\bm{x})\sim\mathcal{P}_{\mathrm{te}}}[(y-\bm{x}_{\mathrm{te}}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|D_{\mathrm{tr}}]\right]$ $\approx\mathcal{B}^{2}(p_{\mathrm{te}},p,a_{1},a_{0})+\mathcal{V}(\sigma^{2},p,p_{\mathrm{tr}},n,d)+\sigma^{2}.$ (37)